### v1

@@ 3.4 Datasets
This section reviews egocentric datasets relevant to egocentric procedural error detection, as shown in Table 2, with a focus on those involving procedural activities. We summarize and analyze these datasets in terms of their scale and data types.
EPIC-Kitchens [damen2020epic] is a large-scale egocentric dataset of 32 participants cooking in diverse home kitchens, totaling 55 hours of unscripted video with audio. The dataset includes rich annotations spanning action categories, temporal boundaries, and participant metadata.
EPIC-Tent [jang2019epic] features outdoor tent-pitching videos from 24 users wearing head-mounted cameras, addressing non-rigid object manipulation scenarios. It uniquely includes participant-provided subjective uncertainty scores.
EgoCom [northcutt2020egocom] comprises 38.5 hours of synchronized binaural audio and egocentric video, with 20,000 timestamped transcriptions and speaker labels from 34 speakers, providing interactive visual and conversational data, including implicit cues absent in third-person datasets.
Assembly 101 [sener2022assembly101] contains 4,321 videos of users assembling and disassembling 101 types of toy vehicles, allowing diverse action sequences, step errors, and corrections. It includes annotations for over 100k coarse- and 1M fine-grained action segments, as well as 18M 3D hand poses.
Ego4D [grauman2022ego4d] comprises over 3,670 hours of daily activity videos from 74 locations in 9 countries, captured with seven types of head-mounted cameras. It includes diverse environments and provides multimodal data, including 3D scans, audio, gaze data, and stereo recordings.
HoloAssist [wang2023holoassist] is tailored for AI assistant training in physical tasks, featuring 166 hours of data from 350 performer-instructor pairs across 20 object manipulation categories. It provides 414 coarse- and 1,887 fine-grained action classes, with manual annotations for actions, error types, and dialogues. Timely error corrections and proactive interventions are included.
Ego-Exo4D [grauman2024ego] features over 1,400 hours of paired egocentric and third-person videos from nearly 800 skilled participants across diverse scenarios, including sports, music, dance, and bicycle repair, capturing both egocentric and contextual views
IKEA Manuals at Work [liu2024ikea] targets furniture assembly with complex scenes, aligning manual instructions and videos for 4D localization of 3D assembly processes. It includes 3D models for 36 items, 98 real assembly videos, and detailed step-level spatiotemporal annotations.
EgoOops [haneji2024egooops] is dedicated to error detection, spanning diverse tasks such as circuit assembly, chemistry experiments, color mixing, block construction, and paper crafts. Unlike prior datasets focused on assembly or cooking, EgoOops offers both correct and incorrect demonstrations with detailed fine-grained annotations.
EgoPER [lee2024error] comprises 386 kitchen cooking videos (28 hours), with 213 showing correct and 173 showing erroneous executions. It features a manually constructed recipe task graph covering all correct and incorrect paths for script-based video generation.
IndustReal [schoonbeek2024industreal] features 84 videos (5.8 hours) of 27 participants assembling 3D-printed toy cars with HoloLens 2. It covers both assembly and maintenance, documenting 724 correct and 38 error steps across 38 error types, including procedural and execution errors.
CaptainCook4D [peddi2024captaincook4d] offers 384 kitchen cooking videos (94.5 hours) across 24 recipes, capturing both regular and error executions. It includes 5.3K step-level and 10K fine-grained action annotations.
@@ 3.5 Evaluation
Precision denotes the proportion of true positive error detections among all predicted errors, reflecting the reliability of the model’s error predictions.
Recall represents the ratio of true positive error detections to all actual error instances, indicating the model’s completeness in identifying errors.
Accuracy is the proportion of correctly classified instances, both erroneous and normal, among all samples, reflecting the overall correctness of the model.
@@
Semantic similarity of error explanations is evaluated using BLEU for n-gram overlap, ROUGE-L for longest common subsequence, and CIDEr for TF-IDF weighted similarity.
Omission Accuracy (O-Acc) quantifies the proportion of detected ground-truth omission errors. Omission Intersection over Union (O-IoU) measures the overlap between predicted and ground-truth omission errors:
@@
Error Detection Accuracy (EDA) is the proportion of segments correctly classified as erroneous or normal, indicating overall segment-level performance in error detection.
Area Under the Curve (AUC) quantifies the model’s discriminative ability by integrating the true positive and false positive rates over varying thresholds, and is commonly applied in binary classification.
Edit Distance measures the dissimilarity between predicted and ground-truth step sequences, enabling assessment of omissions and ordering errors.
@@ 4.1 Definition
Method Years Precision MoF IoU F1 Recall mIoU Jaccard Dataset
ULNI [alayrac2016unsupervised] CVPR 2016 76.0 - - 46.0 67.0 - - YouTube
ULAC [kukleva2019unsupervised] CVPR 2019 - 39.0 - 28.3 - - -
ProcNets-NMS [zhou2018towards] AAAI 2018 30.4 - - 33.4 37.1 33.9 47.6 YouCook2
ProcNets-LSTM [zhou2018towards] AAAI 2018 - - - - - 37.0 50.6
ULAC [kukleva2019unsupervised] CVPR 2019 - 32.9 - - - - - 50salads
ULAC [kukleva2019unsupervised] CVPR 2019 - 41.8 - 26.4 - - - Breakfast
SPS-FL [naing2020procedure] BMVC 2020 - - - 45.4 - - -
SPS-LC [naing2020procedure] BMVC 2020 - - - 53.2 - - -
SPS-FL [naing2020procedure] BMVC 2020 39.6 - - 48.1 64.2 - - Inria
SPS-LC [naing2020procedure] BMVC 2020 47.4 - - 56.0 72.8 - -
JointSeqFL [elhamifar2019unsupervised] ICCV 2019 - - - 28.6 - - - ProceL
CnC [bansal2022my] ECCV 2022 20.7 - - 21.6 22.6 - -
STEPs [shah2023steps] ICCV 2023 23.5 - - 24.9 26.7 - -
@@
OPEL [chowdhury2024opel] NeurIPS 2024 33.6 - 17.9 34.9 36.3 - -
RGWOT [mahmood2025procedure] arXiv 2025 42.2 - 29.4 44.3 46.7 - -
CnC [bansal2022my] ECCV 2022 22.8 - - 22.6 22.5 - - CrossTask
STEPs [shah2023steps] ICCV 2023 26.2 - - 25.9 25.8 - -
GPL [bansal2024united] WACV 2024 24.9 - - 24.5 24.1 - -
OPEL [chowdhury2024opel] NeurIPS 2024 35.6 - 16.9 35.1 34.8 - -
RGWOT [mahmood2025procedure] arXiv 2025 40.4 - 26.3 40.4 40.7 - -
CnC [bansal2022my] ECCV 2022 - - 10.7 22.0 - - - EgoProceL
GPL [bansal2024united] WACV 2024 - - 13.9 25.6 - - -
OPEL [chowdhury2024opel] NeurIPS 2024 - - 16.3 32.0 - - -
RGWOT [mahmood2025procedure] arXiv 2025 - - 31.2 46.8 - - -
@@ 4.2 Unsupervised Learning
Bansal et al. (GPL) [bansal2024united] introduced a graph-based framework with K-means clustering for unsupervised key-step discovery using spatiotemporal and semantic relations. Alayrac et al. (ULNI) [alayrac2016unsupervised] performed text clustering to guide discriminative video clustering for key-step identification. Elhamifar et al. (JointSeqFL) [elhamifar2019unsupervised] employed HMMs and optimization to select representative sub-activity sequences. Kukleva et al. (ULAC) [kukleva2019unsupervised] learned frame embeddings with a neural network and clustered them to infer key steps and order. Chowdhury et al. (OPEL) [chowdhury2024opel] used Optimal Transport theory for video alignment and key-step clustering, thereby avoiding the need for manual annotation. Sener et al. (BP-HMM) [sener2015unsupervised] automatically discovered and parsed semantic steps in large video collections, also generating textual descriptions.
Specifically, GPL [bansal2024united] operates in three stages: segmenting videos into clips using I3D ResNet-50 features to form graph nodes; connecting nodes via semantic and temporal edges and filtering backgrounds through hand-object interaction detection; and optimizing node embeddings with Node2Vec [grover2016node2vec]. Key steps are identified by K-means clustering, with sequence order determined by timestamps.
ULNI [alayrac2016unsupervised] first clusters text narrations by extracting verb-direct object pairs and optimizing an IQP via the Frank-Wolfe algorithm, using WordNet [kilgarriff2000wordnet] for semantic similarity. The top K aligned steps are designated key steps. Subsequently, discriminative clustering of video frame features under text constraints temporally localizes these key steps in the videos.
JointSeqFL [elhamifar2019unsupervised] segments videos and extracts concatenated appearance (VGG16 BoW) and motion (HOF) features, reduced by PCA. HMMs generate sub-activity sequences, with a greedy dynamic programming algorithm and the Frank-Wolfe method used to align and extract global key steps across all videos.
ULAC [kukleva2019unsupervised] extracts frame-level features and timestamps, learning continuous temporal embeddings with an MLP. K-means clusters these embeddings into action classes, and the Viterbi algorithm [richard2018neuralnetwork] infers the optimal sequence of action labels.
OPEL [chowdhury2024opel] is primarily modeled as an optimal transport (OT) problem. It takes two video sequences as input and uses a deep encoder (e.g., ResNet-50) to extract frame-level embeddings. The frame alignment problem is formulated as an OT problem aimed at minimizing the transport cost:
$W_{p}^{p}(f,g)=\min_{\mathbf{T}\in U(\alpha,\beta)}\langle\mathbf{T},\mathbf{D}\rangle,$ (11)
where $D$ denotes the frame-to-frame distance matrix and $T$ is the transport matrix. To enhance performance, OPEL incorporates two types of priors as regularization terms. The Optimality Prior is based on the OT transport matrix $T$, which encourages the alignment of frames that are both semantically similar in the embedding space and temporally close. It is modeled using a Laplacian distribution:
$\mathbf{Q}_{o}(i,j)=\frac{1}{2b}e^{-\frac{|d_{o}(i,j)|}{b}},$ (12)
where $d_{o}(i,j)$ represents the distance of a frame to its optimal aligned position. Temporal Prior promotes the alignment of temporally proximate frames and is also modeled via a Laplacian distribution:
$\mathbf{Q}_{t}(i,j)=\frac{1}{2b}e^{-\frac{|d_{t}(i,j)|}{b}},$ (13)
where $d_{t}(i,j)$ denotes the distance of a frame to the temporal diagonal. The two priors are combined using a dynamic weighting strategy:
$\mathbf{Q}(i,j)=\phi\mathbf{Q}_{t}(i,j)+(1-\phi)\mathbf{Q}_{o}(i,j).$ (14)
BP-HMM [sener2015unsupervised] uses cross-video proposal clustering to extract visual atoms and TF-IDF to select language atoms from ASR captions. Frames are represented multimodally and input to a Beta Process Hidden Markov Model [fox2014joint], with Gibbs sampling used for step inference. This approach identifies step presence, assignments, and frame-step associations, and generates concise textual descriptions with a third-order Markov language model [shannon1948mathematical].
@@ 4.3 Weakly-Supervised Learning
Lin et al. (PLDS) [lin2022learning][anthonio2020wikihowtoimprove] apply distant supervision, generating pseudo-labels by aligning ASR transcripts with external text. Zhou et al. (ProcNets) [zhou2018towards] train using only temporal boundaries, reducing annotation cost. Zhukov et al. (CrossTask) [zhukov2019cross] use shared component classifiers and ordered step lists with narrations for cross-task supervision without temporal labels. Naing et al. (SPS) [naing2020procedure] infer and localize unannotated key steps from partial summaries, framing step completion as joint representation learning and greedy search.
Specifically, PLDS [lin2022learning] consists of three stages. In the first stage, the input is segmented into multiple clips from HowTo100M’s long videos [miech2019howto100m], each containing generated ASR text. Step description texts are retrieved from WikiHow, and pseudo-labels for each clip are obtained by comparing text similarity $S$ :
$\mathcal{S}(a_{l},y_{s}^{(t)})=e(a_{l})^{\top}\cdot e(y_{s}^{(t)}),$ (15)
where $e(a_{l})$ and $e(y_{s}^{(t)})$ $\in\mathbb{R}^{d}$ and $d$ is the dimension of the language embedding space. A Softmax yields a step probability per clip. In the second stage, TimeSformer [bertasius2021space] is trained with Distribution Matching, which minimizes the KL divergence between output and pseudo-label distributions. The resulting representations enable downstream tasks such as procedural learning and action anticipation.
ProcNets [zhou2018towards] is an end-to-end model using ResNet-34 [he2016deep] for frame-level spatial features. A bidirectional LSTM encodes features, and an anchor mechanism generates candidate segments. High-scoring segments are selected and ordered to produce the final step sequence with temporal boundaries.
CrossTask [zhukov2019cross] defines a task as a collection of videos with an ordered step list. For each step, classifiers are trained to identify whether a video segment corresponds to that step. Unlike conventional methods, steps are decomposed into components (e.g., verbs, nouns), with classifiers trained for each component:
[…]
This approach enables component sharing across tasks, allowing new steps to be addressed by recombining existing classifiers. Weak supervision is applied, relying only on step order and matching instructional text. Alternating optimization ensures convergence. The method efficiently recognizes steps without explicit filtering or sorting.
SPS [naing2020procedure] takes video segment features (extracted by VGG and LSTM) and partial key-step annotations $\left\{(Y^{(\ell)},\mathcal{S}_{\ell})\right\}_{\ell=1}^{L}$ as input. The set of entire video key-steps $\{Y^{(\ell)}\}_{\ell=1}^{L}$ and the set of unobserved key-steps $\{\mathcal{U}_{\ell}\}_{\ell=1}^{L}$ are defined. A mapping function projects data into feature space, enabling observed and unobserved key steps to represent the video optimally. Joint optimization of encoding, dissimilarity, and autoencoder losses ensures selected key steps can linearly reconstruct the video (similar to sparse coding):
$\mathcal{L}_{\text{enc},\mathcal{M}}(\theta,\Lambda_{1},\ldots,\Lambda_{L})\triangleq\sum_{\ell=1}^{L}d_{\mathcal{M}}(F_{\Lambda_{\ell}}^{(\ell)},F^{(\ell)}),$ (17)
where $d_{\mathcal{M}}(\cdot,\cdot)$ measures the goodness of the transformed key-steps $F_{\Lambda_{\ell}}^{(\ell)}$ for representing the $\ell$-th video, $F^{(\ell)}$, under the subset selection model $\mathcal{M}$. . The dissimilarity loss encourages the key steps to be dispersed in the feature space to avoid redundancy :
$\mathcal{L}_{dis}(\theta,\Lambda_{1},\ldots,\Lambda_{L})\triangleq-\sum_{\ell=1}^{L}\sum_{i\in\Lambda_{\ell}}\sum_{\binom{j\in\Lambda_{\ell}}{j\neq i}}\|f_{i}^{(\ell)}-f_{j}^{(\ell)}\|_{2}^{2}.$ (18)
Autoencoder loss mitigates overfitting and maintains local data structure through LSTM reconstruction. Alternating Minimization is used for optimization, yielding the complete set of key steps.
@@ 4.4 Self-Supervised Learning
Self-supervised learning creates auxiliary tasks to generate supervisory signals from unlabeled data. As a subset of unsupervised learning, it reduces human annotation effort and enhances generalization for large-scale training.
Bansal et al. (CnC) [bansal2022my] use temporal cycle consistency to learn correspondences and generate pseudo-labels for sequence alignment. Mahmood et al. (RGWOT) [mahmood2025procedure] apply FGWOT with contrastive loss to align frame sequences and create pseudo-labels. Shah et al. (STEPS) [shah2023steps] design unimodal temporal and cross-modal contrastive losses for self-supervision.
Specifically, CnC [bansal2022my] operates in three stages: generating pseudo-labels and frame embeddings with Temporal Cycle Consistency and regularization; modeling clustering as a multi-label graph cut to assign key-step labels; and analyzing temporal occurrences to order key steps, accommodating varied sequences across videos.
RGWOT [mahmood2025procedure] first encodes video frames using 3D convolutional layers and global max pooling, producing 128-dimensional embedding vectors for temporal context representation.For two input videos $X$ and $Y$, the embedding are written as $\mathbf{X}=f_{\boldsymbol{\theta}}(X)\in\mathbb{R}^{N\times D}$ and $\mathbf{Y}=f_{\boldsymbol{\theta}}(Y)\in\mathbb{R}^{M\times D}$, where $f_{\boldsymbol{\theta}}$ is the embedding function and $D$ is the length of the embedding vector. In the second stage, RGWOT aligns frames with FGWOT and contrastive regularization, then constructs a graph and uses $\mathcal{\alpha}$-Expansion to assign key-steps. Final ordering is based on key-step frequency analysis.
STEPS [shah2023steps] primarily consists of two stages. In the first stage, representation learning is performed by inputting multi-modal video data. Used pre-trained networks (VGG-19 and ResNet-50) to extract frame-level appearance features $p_{t}^{appearance}$, while motion features $p_{t}^{motion}$ are extracted using RAFT optical flow [teed2020raft] or OpenPose pose estimation [martinez2019openpose]. A bidirectional LSTM is employed to model the temporal relationships in the feature sequences, generating task-adapted frame-level feature representations. Feature learning is then guided by two loss functions, per-modality temporal loss $\mathcal{L}_{T}$ and cross-cue local loss $\mathcal{L}_{CCL}$.Then it trained the model with the weighted loss function:
$\mathcal{L}=\mathcal{L}_{T}+\lambda_{1}\mathcal{L}_{\mathrm{CCL}}.$ (19)
In the second stage, using K-Means to cluster and extract key-steps. A flowchart of all the mentioned methods is shown in Fig. 6.
Table 4: Summarize the representative datasets in egocentric procedural learning, Ego means egocentric view and Exo means third-person view.

### v2

@@ 3.4 Datasets
This section reviews egocentric datasets relevant to egocentric procedural error detection, as shown in Table 2, with a focus on datasets that involve procedural activities. We summarize and analyze these datasets by scale and data type.
EPIC-Kitchens [damen2020epic] is a large-scale egocentric dataset of 32 participants cooking in diverse home kitchens, totaling 55 hours of unscripted video with audio. The dataset includes rich annotations spanning action categories, temporal boundaries, and participant metadata, but excludes segments with errors. EPIC-Tent [jang2019epic] features outdoor tent-pitching videos from 24 users wearing head-mounted cameras, addressing non-rigid object manipulation scenarios. It uniquely includes participant-provided subjective uncertainty scores, and contains both procedural and execution errors. EgoCom [northcutt2020egocom] comprises 38.5 hours of synchronized binaural audio and egocentric video, with 20,000 timestamped transcriptions and speaker labels from 34 speakers, providing interactive visual and conversational data, including implicit cues absent in third-person datasets, but excludes segments with errors. Assembly 101 [sener2022assembly101] contains 4,321 videos of users assembling and disassembling 101 types of toy vehicles, allowing diverse action sequences, step errors, and corrections. It includes annotations for over 100k coarse- and 1M fine-grained action segments, as well as 18M 3D hand poses. Unlike other datasets, it primarily consists of procedural errors. Ego4D [grauman2022ego4d] comprises over 3,670 hours of daily activity videos from 74 locations in 9 countries, captured with seven types of head-mounted cameras. It includes diverse environments and provides multimodal data, including 3D scans, audio, gaze data, and stereo recordings, but excludes segments with errors. HoloAssist [wang2023holoassist] is tailored for AI assistant training in physical tasks, featuring 166 hours of data from 350 performer-instructor pairs across 20 object manipulation categories. It provides 414 coarse- and 1,887 fine-grained action classes, with manual annotations for actions, error types, and dialogues. Timely error corrections and proactive interventions are included, along with both procedural errors and execution errors. Ego-Exo4D [grauman2024ego] features over 1,400 hours of paired egocentric and third-person videos from nearly 800 skilled participants across diverse scenarios, including sports, music, dance, and bicycle repair, capturing both egocentric and contextual views. Simultaneously encompasses both procedural and execution errors. IKEA Manuals at Work [liu2024ikea] targets furniture assembly with complex scenes, aligning manual instructions and videos for 4D localization of 3D assembly processes. It includes 3D models for 36 items, 98 real assembly videos, and detailed step-level spatiotemporal annotations, but excludes segments with errors. EgoOops [haneji2025egooops] is dedicated to error detection, spanning diverse tasks such as circuit assembly, chemistry experiments, color mixing, block construction, and paper crafts. Unlike prior datasets focused on assembly or cooking, EgoOops offers both correct and incorrect demonstrations with detailed fine-grained annotations, and encompasses both procedural and execution errors. EgoPER [lee2024error] comprises 386 kitchen cooking videos (28 hours), with 213 showing correct and 173 showing erroneous executions. It features a manually constructed recipe task graph covering all correct and incorrect paths for script-based video generation, and encompasses both procedural and execution errors. IndustReal [schoonbeek2024industreal] features 84 videos (5.8 hours) of 27 participants assembling 3D-printed toy cars with HoloLens 2. It covers both assembly and maintenance, documenting 724 correct and 38 error steps across 38 error types, including procedural and execution errors. CaptainCook4D [peddi2024captaincook4d] offers 384 kitchen cooking videos (94.5 hours) across 24 recipes, capturing both procedural and execution errors. It includes 5.3K step-level and 10K fine-grained action annotations.
@@ 3.5 Evaluation
Semantic similarity of error explanations is evaluated using BLEU for n-gram overlap, ROUGE-L for longest common subsequence, and CIDEr for TF-IDF weighted similarity. Omission Accuracy (O-Acc) quantifies the proportion of detected ground-truth omission errors. Omission Intersection over Union (O-IoU) measures the overlap between predicted and ground-truth omission errors:
@@
Error Detection Accuracy (EDA) is the proportion of segments correctly classified as erroneous or normal, indicating overall segment-level performance in error detection. Area Under the Curve (AUC) quantifies the model’s discriminative ability by integrating the true positive and false positive rates across varying thresholds and is commonly used in binary classification. Edit Distance measures the dissimilarity between predicted and ground-truth step sequences, enabling assessment of omissions and ordering errors.
@@ 4.1 Definition
Method Years Precision IoU F1 Recall mIoU Jaccard Dataset
ULNI [alayrac2016unsupervised] CVPR 2016 76.0 - 46.0 67.0 - - YouTube
ULAC [kukleva2019unsupervised] CVPR 2019 - - 28.3 - - -
ProcNets-NMS [zhou2018towards] AAAI 2018 30.4 - 33.4 37.1 33.9 47.6 YouCook2
ProcNets-LSTM [zhou2018towards] AAAI 2018 - - - - 37.0 50.6
ULAC [kukleva2019unsupervised] CVPR 2019 - - 26.4 - - - Breakfast
SPS-FL [naing2020procedure] BMVC 2020 - - 45.4 - - -
SPS-LC [naing2020procedure] BMVC 2020 - - 53.2 - - -
SPS-FL [naing2020procedure] BMVC 2020 39.6 - 48.1 64.2 - - Inria
SPS-LC [naing2020procedure] BMVC 2020 47.4 - 56.0 72.8 - -
JointSeqFL [elhamifar2019unsupervised] ICCV 2019 - - 28.6 - - - ProceL
CnC [bansal2022my] ECCV 2022 20.7 - 21.6 22.6 - -
STEPs [shah2023steps] ICCV 2023 23.5 - 24.9 26.7 - -
@@
OPEL [chowdhury2024opel] NeurIPS 2024 33.6 17.9 34.9 36.3 - -
RGWOT [mahmood2025procedure] arXiv 2025 42.2 29.4 44.3 46.7 - -
CnC [bansal2022my] ECCV 2022 22.8 - 22.6 22.5 - - CrossTask
STEPs [shah2023steps] ICCV 2023 26.2 - 25.9 25.8 - -
GPL [bansal2024united] WACV 2024 24.9 - 24.5 24.1 - -
OPEL [chowdhury2024opel] NeurIPS 2024 35.6 16.9 35.1 34.8 - -
RGWOT [mahmood2025procedure] arXiv 2025 40.4 26.3 40.4 40.7 - -
CnC [bansal2022my] ECCV 2022 - 10.7 22.0 - - - EgoProceL
GPL [bansal2024united] WACV 2024 - 13.9 25.6 - - -
OPEL [chowdhury2024opel] NeurIPS 2024 - 16.3 32.0 - - -
RGWOT [mahmood2025procedure] arXiv 2025 - 31.2 46.8 - - -
@@ 4.2 Unsupervised Learning
Bansal et al. (GPL) [bansal2024united] introduce a graph-based framework with K-means clustering for unsupervised key-step discovery using spatiotemporal and semantic relations. And Elhamifar et al. (JointSeqFL) [elhamifar2019unsupervised] employ HMMs and optimization to select representative sub-activity sequences. Both can effectively filter background frames, handle repeated key-steps, as well as missing or additional actions. However, they rely on the semantic consistency of similar objects and actions, resulting in limited performance for detecting steps involving highly diverse objects or lacking hand-object interactions. In addition, its time complexity grows exponentially with the number of videos. Sener et al. (BP-HMM) [sener2015unsupervised] automatically discover and parse semantic steps in large video collections, generating textual descriptions, and also face the same limitations. Alayrac et al. (ULNI) [alayrac2016unsupervised] performed text clustering to guide discriminative video clustering for key-step identification. It can exploit the complementarity of the two modalities to resolve ambiguities arising from a single modality, and demonstrates strong performance on tasks with highly discriminative textual descriptions, such as ”changing a tire.” To further enable unsupervised learning for activities with unknown categories, Kukleva et al. (ULAC) [kukleva2019unsupervised] learn frame embeddings with an MLP and order key-steps based on the Viterbi algorithm, but it relies excessively on the relative temporal positions of frames, resulting in poor adaptability to action sequences with non-monotonic temporal structures. Chowdhury et al. (OPEL) [chowdhury2024opel] use Optimal Transport theory for video alignment and key-step clustering, thereby avoiding the need for manual annotation, however, it assumes that the objects involved are similar when executing the same key step; large object discrepancies can therefore lead to alignment errors.
@@ 4.3 Weakly-Supervised Learning
Lin et al. (PLDS) [lin2022learning][anthonio2020wikihowtoimprove] apply distant supervision, generating pseudo-labels by aligning ASR transcripts with external text. Still, it relies heavily on a high-quality textual knowledge base; consequently, transcription noise in speech inputs and insufficient coverage of the knowledge base can adversely affect its performance. To reduce annotation cost, Zhou et al. (ProcNets) [zhou2018towards] train using only temporal boundaries, Zhukov et al. (CrossTask) [zhukov2019cross] use shared component classifiers and ordered step lists with narrations for cross-task supervision without temporal labels, and it can further parse previously unseen tasks. Also, Naing et al. (SPS) [naing2020procedure] infer and localize unannotated key steps from partial summaries, framing step completion as joint representation learning and greedy search. Peirone et al. (HiERO) [peirone2025hiero] use a hierarchical graph structure, where a temporal encoder aggregates local temporal information and a function-aware decoder clusters functionally similar segments, and training with video-narration alignment and functional thread losses, the model learns features without task-specific training, achieves strong performance on egocentric video-text alignment and zero-shot procedural learning tasks.
@@ 4.4 Self-Supervised Learning
Shown in Fig. 5, Self-supervised learning creates auxiliary tasks to generate supervisory signals from unlabeled data. As a subset of unsupervised learning, it reduces human annotation effort and enhances generalization for large-scale training.
Bansal et al. (CnC) [bansal2022my] use temporal cycle consistency to learn correspondences and generate pseudo-labels, but rely on the temporal correspondence of key-steps across videos, making it vulnerable to repeated and redundant frames, also ineffective for non-monotonic sequences. To address this issue, Mahmood et al. (RGWOT) [mahmood2025procedure] apply FGWOT with a contrastive loss to align frame sequences and incorporate temporal and structural priors, achieving better performance. Further, designing unimodal temporal and cross-modal contrastive losses, Shah et al. (STEPS) [shah2023steps] do not require video alignment or backbone fine-tuning, support multi-sensor modalities on AR devices, and maintain stable performance in few-shot scenarios.
Table 4: Summarize the representative datasets in egocentric procedural learning, Ego means egocentric view, and Exo means third-person view.